OpenStudy (anonymous):

Segment AB cuts circle Q, what has an area of 36π, in half. If A and B are both points on circle Q, what is the length of segment AB? 2 4 6 8 12

OpenStudy (ghazi):

since Ab cuts in half therefore it must be diameter and diameter= 2* radius pie* r^2=36* pie r=6 and diameter= 6*2

OpenStudy (anonymous):

12

OpenStudy (ghazi):

yea